OpenStudy (anonymous):

Write the matrix corresponding to the system of linear equations 2/3x-3y=4 y=-5

OpenStudy (anonymous):

A X = Y What is A?

OpenStudy (anonymous):

\[ A= \left( \begin{array}{cc} \frac{2}{3} & -3 \\ 0 & -5 \\ \end{array} \right) \] is the matrix of coefficients 2/3 x -3 y 0 x -5 y

OpenStudy (anonymous):

What is X?

OpenStudy (anonymous):

what is X?

OpenStudy (anonymous):

\[ X= \left( \begin{array}{c} x \\ y \\ \end{array} \right) \]

OpenStudy (anonymous):

\[ Y=\left( \begin{array}{c} 4 \\ -5 \\ \end{array} \right) \]

OpenStudy (anonymous):

Do you understand it?

OpenStudy (anonymous):

\[ AX = \begin{pmatrix} \frac{2}{3}x&+&-3y\\ 0x&+&1y \end{pmatrix} \]

OpenStudy (anonymous):

\[ \left( \begin{array}{cc} \frac{2}{3} & -3 \\ 0 & -5 \\ \end{array} \right) \left( \begin{array}{c} x \\ y \\ \end{array} \right)=\left( \begin{array}{c} 4 \\ -5 \\ \end{array} \right) \]

OpenStudy (anonymous):

My last post is the final answer

OpenStudy (anonymous):

\[ AX = \begin{pmatrix} \frac{2}{3}x&+&-3y\\ 0x&+&1y \end{pmatrix} = \begin{pmatrix} \frac{2}{3}&-3\\ 0&1 \end{pmatrix} \begin{pmatrix} x\\ y \end{pmatrix} \]

OpenStudy (anonymous):

I guess this would be the most thorough way to show it:\[ AX = \begin{pmatrix} \frac{2}{3}x-3y\\ y \end{pmatrix} = \begin{pmatrix} \frac{2}{3}x&+&-3y\\ 0x&+&1y \end{pmatrix} = \begin{pmatrix} \frac{2}{3}&-3\\ 0&1 \end{pmatrix} \begin{pmatrix} x\\ y \end{pmatrix} \]And \(Y\) is very simple: \[ Y = \begin{pmatrix} 4\\ -5 \end{pmatrix} \]